Yuki (age 45 at year-end) has been contributing to a traditional IRA for years (all deductible contributions), and her IRA is now worth $39,000. She is trying to decide whether she should roll over her traditional IRA into a Roth IRA. Her current marginal tax rate is 24 percent. She plans to withdraw the entire balance of the account in 20 years and she expects to earn a before-tax rate of return of 4.9 percent on her retirement accounts and a 3.9 percent after-tax rate of return on all investments outside of her retirement accounts. For each of the following alternative scenarios, indicate how much more or less Yuki will accumulate after taxes in 20 years if she rolls over her traditional IRA into a Roth IRA. Be sure to include the opportunity cost of having to pay taxes on the rollover. (Use Table 1.) (Round your intermediate calculations and final answers to the nearest whole dollar amount.)

a. When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 35 percent.

    

Accumulation if she keeps funds in traditional IRA______

Accumulation if she rolls funds into Roth IRA___________

Accumulation if she rolls over traditional IRA into Roth IRA______

b. When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 18 percent.

    

Accumulation if she keeps funds in traditional IRA
Accumulation if she rolls funds into Roth IRA
Lessselected answer correct accumulation if she rolls over traditional IRA into Roth

c. Assume the same facts as in part (b), except that she earns a 2.9 percent after-tax rate of return on investments outside of the retirement accounts.

    

Answer is not complete.

Accumulation if she keeps funds in traditional IRA
Accumulation if she rolls funds into Roth IRA
Greaterselected answer correct accumulation if she rolls over traditional IRA into Roth IRA   
Answer #1

ANSWER

Part A

Accumulation if she keeps funds in traditional IRA = 39,000*(1.049^20) *(1-35%) = $65,991

Accumulation if she rolls funds into Roth IRA = (39,000*(1.049^20))-(39,000*24%*(1.039^20)) = $81,407

Difference = 81,407-65,991= $15,416

Therefore,

Accumulation if she keeps funds in traditional IRA

$65,991

Accumulation if she rolls funds into Roth IRA

$81,407

Greater

Accumulation if she rolls over traditional IRA into Roth IRA

$15,416

Part B

Accumulation if she keeps funds in traditional IRA = 39,000*(1.049^20) *(1-18%) = $83,251

Accumulation if she rolls funds into Roth IRA = (39,000*(1.049^20))-(39,000*24%*(1.039^20)) = $81,407

Difference = 81,407-83,251= $1,844

Therefore,

Accumulation if she keeps funds in traditional IRA

$83,251

Accumulation if she rolls funds into Roth IRA

$81,407

Lesser

Accumulation if she rolls over traditional IRA into Roth IRA

$1,844

Part C

Accumulation if she keeps funds in traditional IRA = 39,000*(1.049^20) *(1-18%) = $83,251

Accumulation if she rolls funds into Roth IRA = (39,000*(1.049^20))-(39,000*24%*(1.029^20)) = $84,945

Difference = 84,945-83,251= $1,694

Therefore,

Accumulation if she keeps funds in traditional IRA

$83,251

Accumulation if she rolls funds into Roth IRA

$84,945

Greater

Accumulation if she rolls over traditional IRA into Roth IRA

$1,694

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************